OpenStudy (anonymous):

describe the general solution in degrees: 6tan(piθ)=36

OpenStudy (anonymous):

this is as far as I got: tan(piθ) = 6, not sure what to do after that...the pi in the () is messing me up.

OpenStudy (anonymous):

\[\pi\theta=\arctan(6)\] \[\theta=\frac{\arctan(6)}{\pi}\]